OpenStudy (anonymous):

a golf ball is hit off the ground at an angle of pi/6 and travels 400 feet. How long was it in the air???

OpenStudy (anonymous):

V = 400/(t*cos(pi/6)) (g/(2*sin(pi/6)))*t = V (g/(2*sin(pi/6)))*t =400/(t*cos(pi/6))

OpenStudy (anonymous):

t^2 = (2*400*tan(pi/6))/g

OpenStudy (anonymous):

edited: left out a parenthesis.)